### v1

# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

Abstract
Despite the outstanding performance of current deep learning models across various domains, their fundamental bottleneck in abstract reasoning remains unresolved. To address this challenge, the academic community has introduced Raven’s Progressive Matrices (RPM) problems as an authoritative benchmark for evaluating the abstract reasoning capabilities of deep learning algorithms, with a focus on core intelligence dimensions such as abstract reasoning, pattern recognition, and complex problem-solving. Therefore, this paper centers on solving RPM problems, aiming to contribute to enhancing the abstract reasoning abilities of machine intelligence. Firstly, this paper adopts a “causal chain modeling” perspective to systematically analyze the complete causal chain in RPM tasks: image $\rightarrow$ abstract attributes $\rightarrow$ progressive attribute patterns $\rightarrow$ pattern consistency $\rightarrow$ correct answer. Based on this analysis, the network architecture of the baseline model DIO is designed. However, experiments reveal that the optimization objective formulated for DIO, namely maximizing the variational lower bound of mutual information between the context and the correct option, fails to enable the model to genuinely acquire the predefined human reasoning logic. This is attributed to two main reasons: the tightness of the lower bound significantly impacts the effectiveness of mutual information maximization, and mutual information, as a statistical measure, does not capture the causal relationship between subjects and objects. To overcome these limitations, this paper progressively proposes three improvement methods:
  1. 1.
    Brando: By generating trainable hypothetical incorrect options, the Brando explicitly tightens the variational lower bound of mutual information.
  2. 2.
    WORLD: Abandoning the generative assumption, the WORLD constructs a Gaussian mixture model for the feature distribution of RPM images, providing an infinite and informative set of incorrect options through weighted sampling to further tighten the lower bound.
  3. 3.
    DIEGO: Introducing metadata supervision, DIEGO directly corrects semantic biases in the “abstract attributes $\rightarrow$ progressive patterns” link of the causal chain, aligning the model’s internal representations with human reasoning frameworks.
Experiments demonstrate that these three methods not only significantly enhance the accuracy of discriminative RPM tasks but also empower DIO to solve open-ended generative RPM problems. The findings of this paper offer causal-driven architectural design principles, learning objective refinement strategies, and novel insights into cross-modal generation for abstract reasoning research.
Index Terms:
Abstract Reasoning, Raven’s Progressive Matrices, Generative RPM, Variational Mutual Information, Causal Chain Modeling, Representation Learning.
I Introduction
Deep learning, a cross-disciplinary technology that draws inspiration from the operational mechanisms of the human brain, has in recent years showcased formidable potential and transformative capabilities across numerous fields, with particularly outstanding accomplishments in generative tasks. Key technologies, which include Generative Adversarial Networks (GANs)[1], Variational Autoencoders (VAEs)[2], Diffusion Model[3], and the Transformer architecture that is based on the attention mechanism[4], empower models to generate high-quality and realistic images, texts, and videos from random noise or given conditions, thereby expanding the application scope of artificial intelligence in the realm of content creation.
In computer vision, deep learning, which makes use of deep convolutional neural networks and their variants, can efficiently extract and accurately classify image features, outperforming traditional methods in tasks such as object detection and image segmentation [5, 6, 7]. In the field of natural language processing, language models that are built on deep learning can automatically learn language grammar, semantics, and contextual associations, performing exceptionally well in tasks such as machine translation and text generation, and propelling the technology towards intelligence and automation [4, 8, 9]. In the multimodal cross-field of visual question answering systems, deep learning, which integrates visual and language information, can achieve joint understanding and reasoning, accurately answering natural language questions that are related to images [11, 12].
Currently, research on Large Language Models (LLMs) is a focal point in the field of deep learning. These models, which possess a large number of parameters and are trained on massive data, demonstrate strong generalization ability in various natural language processing tasks, providing efficient and unified solutions for different tasks and facilitating the progress of artificial intelligence towards generalization and intelligence. However, the academic community has recently found that although LLMs have achieved remarkable results in language understanding and generation, they encounter obvious performance bottlenecks in abstract reasoning tasks[13, 14, 15, 16, 17]. Abstract reasoning ability, which is a vital component of human advanced intelligence and involves the understanding of abstract concepts, the inference of logical relationships, and the formulation of problem-solving strategies, serves as a key indicator for measuring the intelligence level of models. The shortcomings of LLMs not only expose the limitations of existing models in handling complex cognitive tasks but also underscore the significance of enhancing their abstract reasoning ability for achieving more powerful and general artificial intelligence.
Based on the above, this paper centers on the issue of improving the abstract reasoning ability of deep learning models. It explores effective model architecture design, training strategy optimization, and data augmentation methods, which aim to offer theoretical and practical support for breaking through the performance bottlenecks of existing models in abstract reasoning tasks and promoting the application and expansion of deep learning technology in a broader range of cognitive tasks.
II Problem Description
In artificial intelligence, abstract reasoning ability stands as a core hallmark of intelligence and a critical prerequisite for achieving human-level or superhuman AI. Yet current AI technologies are constrained by heavy reliance on large-scale annotated data, performing poorly in abstract tasks requiring causal reasoning, such as medical diagnosis and financial analysis. While academic efforts have produced datasets like RAVEN [19] and PGM [20] for benchmarking reasoning capabilities, designing algorithms to handle symbolic logic, causality, and uncertainty remains challenging. Strengthening this ability demands algorithmic innovation integrated with interdisciplinary insights from cognitive science, neuroscience, and beyond, enabling breakthroughs through cross-domain collaboration.
II-A Raven’s Progressive Matrices
Raven’s Progressive Matrices (RPM) [18] serves as an authoritative benchmark for evaluating the advanced cognitive capabilities of deep learning algorithms, focusing on core intelligence dimensions such as abstract reasoning, pattern recognition, and complex problem-solving. Its standardized testing framework demonstrates high correlation with human intelligence metrics, making it a reliable and stable tool for assessing the cognitive complexity of machine learning models. In current research, the RAVEN dataset [19] and PGM dataset [20] have emerged as the two most representative standardized test paradigms in the RPM task domain.
The RAVEN dataset comprises structured RPM instances, with each test unit containing 16 images: 8 forming a “context matrix” as the reasoning basis, and the remaining 8 constituting a candidate option pool. The task requires selecting appropriate images from the options to construct a complete $\times3$ matrix that adheres to progressive geometric transformation rules and abstract logical principles. Figure 1 (left) illustrates a typical instance from the RAVEN dataset. Similarly, while the PGM dataset also employs 8 images for context and options, its underlying logical rules must satisfy dual constraints in both row and column directions of the matrix. Figure 1 (right) presents a standard instance from the PGM dataset.
Figure 1: RAVEN and PGM Instance
II-B Related work
Image reasoning models such as CoPINet[23], LEN+teacher[22], and DCNet[24] primarily focus on learning differences and rules. Meanwhile, models like NCD[26], SAVIR-T[27], as well as neuro-symbolic systems (including PrAE, NVSA, ALANS[28, 29, 30]), have achieved promising results in both enhancing model interpretability and improving reasoning accuracy. RS-CNN and RS-TRAN[31] demonstrate exceptional capabilities in addressing problems related to RPM, while Triple-CFN[33] stands out among numerous models by implicitly extracting and indexing information about concepts and reasoning units, effectively enhancing reasoning accuracy.
III Methodology
This paper first proposes that to enhance the ability of deep learning models to solve RPM, which are causal reasoning problems, when designing the model’s structure and dividing its modules, one should fully refer to the causal chains within the problems. Based on this methodology, this paper designs a new model for solving RPM problems, named DIO. However, this paper finds that DIO is in satin without substance. Its learning approach is essentially the same as that of the previous RPM solving model, whose purpose is to maximize the mutual information between the context matrix and the correct option in RPM instance. This learning approach leads to a situation where, no matter how well its structure aligns with the causal chains in the problems, its learning and optimization will not progress towards the human predefined reasoning framework. Therefore, this paper proposes three methods, namely Brando, WORLD, and DIEGO, to refine DIO’s learning objectives. As a bonus, when DIO adopts DIEGO method, it demonstrates the capability to solve generative RPM problems.
The images in the RPM instances of this paper have adopted the annotation approach as illustrated in Figure 2, which enables clearer explanation of our model and method.
Figure 2: Annotations of Images Within an RPM Instance.
Spcifically, we denote the 16 RPM images as $\{x_{i}|i\in[1,16]\}$.
IV DIO: A Baseline Model
In this section, we design a high-performance RPM solving model, named DIO (Deep-learning Intelligent-model Organized via Causal Chains). While providing a detailed exposition of the model’s structure, we have also elucidated its underlying design principles.
IV-A The Causal Chains in RPM Problems.
The solving process of RPM problems is as follows: The solving model needs to discern the progressive pattern of abstract image attributes from the context matrices $\{x_{i}|i\in[1,8]\}$ in an RPM instance. Subsequently, it should select the ideal option $x_{\alpha}$ from the eight options $\{x_{i}|i\in[9,16]\}$ in the option pool that will not disrupt this progressive pattern [35]. This implies that there exists a causal chain in RPM problems: RPM images $\rightarrow$ abstract attributes $\rightarrow$ progressive patterns of attributes $\rightarrow$ consistency of progressive patterns $\rightarrow$ correctness of options.
This paper proposes that if the structural design of the RPM solving model is guided by this causal chain, the constructed model will demonstrate high performance. This prompted us to design the DIO model. In accordance with the causal chains in RPM problems, the DIO model is divided into four modules: the image feature extraction module, the progressive pattern induction module, the progressive pattern consistency evaluation module, and the option evaluation module.
IV-B The Image Feature Extraction Module.
RPM problems place significant emphasis on evaluating the abstract reasoning capabilities of deep models. Consequently, when extracting features from RPM images, we opt for a Visual Transformer (ViT) [21] that prioritizes image context. Figure 3 illustrates the process by which ViT extracts tokenized image features $\{z_{ij}|j\in[1,N]\}$ from the input image $x_{i}$, where $j$ denotes the token index of the image features and $N$ denotes the total patch count.
Figure 3: The Image Feature Extraction Module.
This process can be formally expressed as:
$\{z_{ij}|i\in[1,16],j\in[1,N]\}=\text{FEM}(x_{i})|_{i=1}^{16}$ (1)
IV-C The Progressive Pattern Induction Module
The deliberate design of RPM problems necessitates that participants must simultaneously observe two rows and two columns to identify the progressive patterns within the matrix. This represents a critical link in the causal chain of RPM problems. Consequently, the progressive pattern induction module of DIO adopts the structure illustrated in Figure 4.
Figure 4: The Progressive Pattern Induction Module.
Figure 4 indicates that after incorporating the features of option $\{z_{\alpha j}|j\in[1,N]\}$ into the context matrix $\{z_{ij}|i\in[1,8],j\in[1,N]\}$, DIO will summarize the progressive patterns among the features that share the same token index $j$. DIO first utilizes an MLP (Multi-Layer Perceptron) to extract row and column features from the matrix. Subsequently, it randomly selects two row features from the three available row features and two column features from the three column features, combining these four features for further processing. After binding four trainable vectors of the same dimension to the selected four features, DIO employs a Transformer Encoder to process them into eight logical outputs. Among the output vector groups, the four vectors corresponding to the trainable vectors are denoted as $\{P_{jl}|l\in[1,4]\}$.
$\{P_{jl}|l\in[1,4]\}$ is regarded as the partial representation of progressive patterns within the matrix. If all the $C_{3}^{2}\cdot C_{3}^{2}=9$ possible ways of selecting row and column features are traversed (where $C_{n}^{r}$ denotes the number of combinations of $n$ items taken $r$ at a time), then a comprehensive representation of progressive patterns, $\{P_{jkl}|k\in[1,9],l\in[1,4]\}$, can be obtained. This module can be formally expressed as:
$\{P_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$
$=\text{PPIM}(\{z_{ij},z_{\alpha j}|i\in[1,8]\})|_{j=1}^{N}$ (2)
IV-D The Progressive Pattern Consistency Evaluation Module
As its name suggests, DIO’s progressive pattern consistency evaluation module is responsible for using a Transformer Encoder to compute the consistency among progressive pattern representations that share the same $j$-index and $l$-index. Specifically, DIO applies a Transformer Encoder to perform dimension-invariant processing on $\{P_{\alpha jkl}|k\in[1,9]\}$, resulting in $\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$. Subsequently, each vector in $\{P^{\prime}_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$ is mapped to a scalar via an MLP, serving as a logical score $s_{\alpha jkl}$. This process can be formally expressed as:
$\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$ $=\text{Trans-Encoder}(\{P_{\alpha jkl}|k\in[1,9]\})$ (3)
$s_{\alpha jkl}$ $=\text{MLP}(P^{\prime}_{\alpha jkl})$ (4)
The entire module can be represented as:
$\{s_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$
$=\text{PCEM}(\{P_{\alpha jkl}|k\in[1,9]\}_{i})|_{j=1,l=1}^{N,4}$ (5)
IV-E The Option Evaluation Module.
The DIO’s option evaluation module computes the score $S_{\alpha}$ that reflects the correctness of option $x_{\alpha}$ by summing $\{s_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$ according to the formula:
$S_{\alpha}$ $=\frac{1}{N\cdot 9\cdot 4}\big{(}\sum_{j=1}^{N}\sum_{k=1}^{9}\sum_{l=1}^{4}s_{\alpha jkl}\big{)}$ (6)
Use the softmax function to map $\{S_{i}|i\in[8,16]\}$ to a probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$:
$\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{S_{\alpha}}}{\sum_{{\hat{\alpha}}=9}^{16}e^{S_{{\hat{\alpha}}}}}$ (7)
We introduce a constraint mechanism that correlates the DIO-modeled probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$. This constraint enables joint optimization of all trainable parameters across DIO’s modules:
${\ell}_{\text{DIO}}=-\mathbb{E}\left[\log\frac{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\sum_{\hat{\alpha}=9,\hat{\alpha}\neq\alpha}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (8)
Where $\mathbb{E}(\cdot)$ denotes the expectation operation.
V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM Solving Model
This paper proposes that architecture optimization based on causal chains can effectively enhance the performance of deep learning networks in solving RPM problems. However, through further research, we have found that this methodology still cannot fundamentally improve the logical reasoning capabilities of deep learning networks.
V-A The learning objective of the DIO
By abstracting away the specific architecture of DIO (i.e., merging equations (1)-(6)), we obtain a unified formulation that highlights its core inductive mechanism:
$S_{\alpha}$ $=\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})$ (9)
Leveraging the probabilistic framework in Equation (7), DIO can be interpreted as a tunable probabilistic model where $\theta$ represents learnable parameters:
$\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (10)
Notably, the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ can be expressed as:
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}P(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (11)
Where $K$ denotes the number of incorrect options. If we consider $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a sufficient approximation of the true conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$ , then substituting the $\hat{P}$ defined in Equation (10) into the Equation (12) yields a mutual information lower bound expressed by DIO:
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]$ (12)
When $K\geq 1$, that is, when there exists at least one incorrect option, the lower bound aligns with DIO’s optimization objective, as expressed by the following equation:
$\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]=-\ell_{\text{DIO}}+\log K$ (13)
Equation (13) reveals that DIO’s learning objective (i.e. maximizing $\ell_{\text{DIO}}$) is neither the human-like reasoning logic embedded in RPM problems nor even the suboptimal approach of directly maximizing mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, but rather optimizing the variational lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. Furthermore, based on the properties of logarithmic functions, the following inequality holds:
${\ell}_{\text{DIO}}$ $<\mathbb{E}\left[\log\frac{\sum_{\hat{\alpha}=9}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8})}\right]$
$=-1\cdot\log\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ (14)
The right-hand side of the Inequality (V-A) embodies the classical formulation for optimizing the probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the cross-entropy (CE) loss function, which is the typical paradigm adopted by most previous end-to-end RPM solving models. From this, it can be inferred that for those end-to-end RPM solving models that have also focused on constructing the probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ in the past, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
Equation (13) and Inequality (V-A) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In this case, the looseness of the lower bound undoubtedly has a direct and significant impact on the performance of these models.
V-B Introducing Hypothetical Options to Derive a Tight Lower Bound for Mutual Information.
This paper argues that employing DIO to derive a tighter variational lower bound and directing its optimization objective toward the maximization of this new lower bound can more effectively maximize mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, thereby further DIO’s performance.
By analyzing Formula (12), we deduce that increasing the number of incorrect options $K$ encountered by DIO can guide its learning objective toward a tighter variational lower bound. However, RPM problem instances only offer a limited number (seven) of incorrect options. To this end, this paper attempts to introduce additional hypothetical incorrect options for DIO. Specifically, we introduce a set of hypothetical options $\{x_{i}|i\in[17,\beta]\}$ into the training pipeline of DIO. Consequently, the conditional distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ is reparameterized as:
$\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)=\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (15)
Correspondingly, the optimization objective of DIO, the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, is updated as follows:
$I\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{\beta-9}\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]$ (16)
Intuitively, the looseness of the updated lower bound has been reduced.
However, In Formula (16), it is not difficult to observe that if these hypothetical options cannot change DIO’s decision-making then this new lower bound, when serving as an optimization objective, will not undergo any fundamental improvement. In other words, when the probabilities assigned by the DIO model to the hypothetical options $\{x_{i}\mid i\in[17,\beta]\}$ approach zero, which means that these options do not alter the probability distribution $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16]\}$ that DIO assigns to the original options $\{x_{i}|i\in[9,16]\}$, the variational lower bounds derived from Equation (16) and Equation (12) become equivalent given that both function as loss functions for neural networks. This implies that unconstructive options contribute nothing to tightening the lower bound.
Therefore, this paper proposes introducing extra constructive hypothetical options into DIO—options that reinforce its endorsement of the correct options while simultaneously strengthening its repulsion of the incorrect ones—so as to steer DIO’s learning objective toward a tighter variational lower bound and ultimately further enhance its performance.
VI Brando: Tightening the Variational Lower Bound of Mutual Information by Constructing Hypothetical Options.
This paper has put forward the idea that incorporating extra constructive hypothetical options into DIO’s training process offers a promising avenue for boosting its performance. To this end, this section introduces a training framework specifically designed to furnish DIO with such hypothetical options, thereby enhancing its overall effectiveness.
VI-A Trainable Hypothetical Opions.
The RPM problem expects the training process of the solving model to involve only a limited number of supervised signals, such as the index of correct options. Relying on manual design to create constructive options goes against the original intent of RPM problems. Therefore, this paper proposes training a set of constructive options instead. This methodology is referred to as Brando (Bounded Random Options Mapping Method) in this paper.
This paper establishes a set of trainable hypothetical options $\{x_{i}|i\in[17,\beta]\}$ for the context matrix $\{x_{i}|i\in[1,8]\}$. This paper is not concerned with the specific attributes of these options; rather, it focuses solely on how much these options can alter DIO’s evaluation of the original options $\{x_{i}|i\in[9,16]\}$. In other words, we are solely concerned with the constructiveness of these hypothetical options for DIO, rather than their specific configurations.
Since these options are expected to improve $S_{\alpha}$ while reducing $\{S_{\hat{\alpha}}|{\hat{\alpha}}\in[9,16],{\hat{\alpha}}\neq\alpha\}$, this paper designs a new loss function $\ell_{\text{Brando}}$ to train these hypothetical options and further optimize DIO:
$\ell_{\text{Brando}}=\ell_{\text{up}}+\ell_{down}$ (17)
Where $\ell_{\text{up}}$:
$\ell_{\text{up}}=-\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{\beta}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (18)
and $\ell_{\text{down}}$:
$\ell_{\text{down}}=-\frac{1}{\beta-17}\sum_{{\hat{\beta}}=17}^{\beta}\log\frac{e^{\text{DIO}(\{x_{i},x_{\hat{\beta}}|i\in[1,8]\})}}{\sum_{{\hat{\alpha}}=9,{\hat{\alpha}}\neq\alpha}^{16}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}$ (19)
In these noval loss function:
  1. 1.
    It can be observed that $\ell_{\text{up}}$ requires the probability $\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ derived by DIO to be greater than all elements in the set $\{\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|{\hat{\alpha}}\in[9,\beta],{\hat{\alpha}}\neq\alpha\}$. In other words, $\ell_{\text{up}}$ demands that DIO be able to identify the correct option amidst the interference from the hypothetical options $\{x_{i}|i\in[17,\beta]\}$.
  2. 2.
    On the other hand, $\ell_{\text{down}}$ constrains that any element in the set $\{\hat{P}(x_{\hat{\beta}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\beta}\in[17,\beta]\}$ must be higher than all elements in the set $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16],\hat{\alpha}\neq\alpha\}$. In other words, $\ell_{\text{down}}$ ensures hypothetical options $\{x_{i}\mid i\in[17,\beta]\}$ outscore the original incorrect options $\{x_{i}\mid i\in[9,16],i\neq\alpha\}$ under DIO.
In loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive options. When optimizing the DIO using $\ell_{\text{Brando}}$ as the loss function, it can induce bounded hypothetical options while leveraging a tighter variational lower bound to enable DIO to maximize the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.
VI-B Map Out Hypothetical Options
Formulas (18) and (19) indicate that hypothetical options should demonstrate constructiveness for any given context matrix, meaning that hypothetical options should be customized according to the given context matrix. Therefore, training a set of shared hypothetical options is inappropriate. This implies that we require a mapping with the following form:
$\{x_{i}|i\in[17,\beta]\}=\mathcal{F}(\{x_{i}|i\in[1,8]\})$ (20)
Evidently, the mapping $\mathcal{F}(\cdot)$ is required to derive a variable number ($\beta$) and diverse hypothetical options for the context matrix $\{x_{i}|i\in[1,8]\}$. This indicates that conventional deep neural networks struggle to undertake the task of the mapping $\mathcal{F}(\cdot)$. Firstly, constructing a network with variable and dynamic output specifications is highly challenging. Additionally, from the perspective of neural network optimization, achieving a many-to-many mapping is infeasible. Therefore, this paper proposes a many-to-one mapping approach:
$x_{\hat{i}}=\hat{\mathcal{F}}(\{x_{i},o_{\hat{i}}|i\in[1,8]\})$ (21)
$\text{Where}\quad o_{\hat{i}}\sim N(\mu,\sigma^{2})$
[…]
$\{x_{\hat{i}}|i\in[17,\beta]\}=\hat{\mathcal{F}}(\{x_{i},o_{\hat{i}}|i\in[1,8]\})\mid_{\hat{i}=17}^{\beta}$ (22)
[…]
$z_{\hat{i}j}=\text{Brando}(\{z_{ij},o_{\hat{i}}|i\in[1,8]\})$ (23)
$\text{Where}\quad o_{\hat{i}}\sim N(\mu,\sigma^{2})$
Formula (23) shows that we move away from the option-mapping paradigm and instead employs a representation-mapping approach, easing the learning pressure on Brando. According to Formula (23), we have designed the structure of the Brando network as shown in the Figure 5.
Figure 5: The Brando network.
As indicated in Figure 5, the Brando network is formed by stacking multiple logical layers, with no parameter sharing between each layer. The input to each logical layer consists of $\{z_{ij}|i\in[1,8]\}$ and $o_{\hat{i}}$. Except for the first layer, the input to all subsequent layers comes from the output of the preceding layer. For the first layer, the $\{z_{ij}|i\in[1,8]\}$ is derived from the representations of the context matrix, while the $o_{\hat{i}}$ is obtained through the reparameterization trick [2] from a learnable Gaussian distribution parameterized by mean and variance $N(\mu,\sigma^{2})$. For the output $o_{\hat{i}}$ from the final logic layer, this paper treats it as the representation $z_{\hat{i}j}$ of a hypothetical option provided by Brando for the context matrix $\{z_{ij}|i\in[1,8]\}$.
For the specific structure of each logical layer, it consists of a set of mapping matrices $\{W_{q},W_{k},W_{v}\}$, a multi-head self-attention mechanism, a feed-forward layer, a memory bank $M_{n}$, a buffer $B_{n}$, and a reparameterization head.
  1. 1.
    When the input set $\{z_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ is fed into a single logical layer, the mapping matrices $W_{q}$, $W_{k}$, and $W_{v}$ transform this input into query ($Q$), key ($K$), and value ($V$) representations, respectively.
  2. 2.
    After these query ($Q$), key ($K$), and value ($V$) representations are appended positional encoding and processed by layer normalization, the multi-head self-attention mechanism computes interactions among them, yielding the attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$.
  3. 3.
    As indicated by the computational step 1 in the diagram, the normalized $Q$ representations employ a linear attention mechanism to learn the historical input representations stored in the memory bank $M_{n-1}$ and buffer $B_{n-1}$, resulting in the output set $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}$. The learning process can be interpreted as:
    $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}=\frac{\phi(Q)M_{n-1}}{\phi(Q)B_{n-1}}$ (24)
    Where basis function $\phi(\cdot)$ is an element-wise ELU+1 operation.
  4. 4.
    In computation steps 2 and 3, the memory bank $M_{n-1}$ and buffer $B_{n-1}$ are updated by leveraging the key ($K$) and value ($V$) mappings derived from the current input representations. The update process can be expressed as:
    $M_{n}$ $=M_{n-1}+\phi(K)^{T}\left(V-\frac{\phi(K)M_{n-1}}{\phi(K)B_{n-1}}\right)$ (25)
    $B_{n}$ $=B_{n-1}+\sum\phi(K)$ (26)
  5. 5.
    After the learnable scalar $\tau$ is compressed into a scaling coefficient via the sigmoid function, a weighted summation is performed on $\tilde{o}_{\hat{i}}$ and $\hat{o}_{\hat{i}}$ using this coefficient to update $\hat{o}_{\hat{i}}$, enabling it to balance its own reasoning with image representations from other instances.
    $\hat{o}_{\hat{i}}\leftarrow[1-\text{Sigmoid}(\tau)]\hat{o}_{\hat{i}}+\text{Sigmoid}(\tau)\tilde{o}_{\hat{i}}$ (27)
  6. 6.
    The attention outputs $\{\hat{z}_{ij}|i\in[1,8]\}$ with the updated $\hat{o}_{\hat{i}}$, which has undergone skip connection with logical inputs, is processed by the feed-forward layer and subjected to another skip connection, yielding the logical output $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}$. This process can be expressed as:
    $Z\leftarrow\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}+\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ (28)
    $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}\leftarrow\text{FFL}(Z)+Z$ (29)
    Here, FFL()($\cdot$) refers to the feed-forward layer.
  7. 7.
    It is noteworthy that only after the $^{o}_{\hat{i}}$ within the logical output $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}$ have been processed by the reparameterization head can they serve as the final output. The process of the reparameterization head can be interpreted as follows:
[…]
    ${o}_{\hat{i}}\leftarrow{o}_{\mu}+e^{\frac{o_{\ln\sigma}}{2}}\cdot\epsilon$ (31)
    The $\epsilon$ represents a random sample drawn from $N(0,1)$ and MLP()($\cdot$) refers to a Multi-Layer Perceptron capable of doubling the dimension of $^{o}_{\hat{i}}$.
It is not hard to find that the memory bank $M_{n}$ and buffer $B_{n}$ in each logical layer will incorporate new memories as a new batch of context matrices is observed. Therefore, the initial form of the memory bank requires a clear definition.
$M_{0}=0,\,B_{0}=1\times 10^{-9}$ (32)
Furthermore, the first update of the memory bank and the buffer can be explained as follows:
$M_{1}=\phi(K)^{T}V,\,B_{1}=\sum\phi(K)$ (33)
From the learnable Gaussian distribution ${N}({\mu},{\sigma}^{2})$, we sample random vectors $^{o}_{\hat{i}}$, which undergo sequential refinement in the Brando network through dynamic interaction with both current context matrix representations and historical image representations accumulated across multi-round training iterations. This iterative conditioning process enables $^{o}_{\hat{i}}$ to progressively assimilate visual patterns observed throughout the training trajectory, ultimately yielding the ultimate hypothesis representation ${z}_{\hat{i}j}$. Moreover, alterations in the initial promoter $^{o}_{\hat{i}}$ are capable of inducing changes in the ultimate option representation $z_{\hat{i}j}$. This satisfies the demand to map diverse hypothetical options for the context matrix.
VI-C The Combination of Brando and DIO.
Prior to integrating the Brando network with DIO, DIO must be pre-trained on the original option pool $\{x_{i}|i\in[9,16]\}$ using the loss function $\ell_{\text{DIO}}$. Once DIO is pre-trained to approach its optimal performance, the Brando network will be integrated with DIO according to the method illustrated in Figure 6.
Figure 6: The Combination Method of DIO and Brando.
Subsequently, the loss function will be switched from $\ell_{\text{DIO}}$ to $\ell_{\text{Brando}}$ for continued training of this integrated system, thereby further optimizing DIO’s performance. Another key point is that during the joint training process, DIO’s parameters excluding those in the image feature extraction module will be frozen every other iteration.
VII WORLD: Tightening the Variational Lower Bound of Mutual Information Through Deconstructing the Problem Space
It is readily apparent that the architecture of the Brando network leaves substantial room for theoretical refinement. The pursuit of a more optimal structure for implementing the mapping $\mathcal{F}(\cdot)$ appears to be an open-ended endeavor, with no discernible asymptote in sight. Moreover, providing a sufficient number of hypothetical options for DIO using Brando entails considerable computational overhead. Therefore, this paper argues that Brando is not a flawless methodology. To this end, this paper proposes a method named WORLD (Weighted Option-Represented Lower-bound Delimitation Method), which breaks away from the paradigm of the mapping $\mathcal{F}(\cdot)$. This method can also tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
VII-A Modeling is Preferable to Estimation.
Formula (12) indicates that the key to guiding the learning objective of solving model towards a tighter lower bound of the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$ lies in providing additional incorrect options to the model for each context matrix. Moreover, this paper also points out that the introduced additional incorrect options must avoid being unconstructive, that is, after being incorporated into the training process of the solving model, they should be able to enhance the decision-making capability of the solving model.
The Brando network essentially functions as a distribution transformation mechanism, whose task is to map samples $o_{\hat{i}}$ drawn from a Gaussian distribution $N(\mu,\sigma^{2})$ into features $z_{\hat{i}j}$ of incorrect options customized for a given context matrix. This paper argues that instead of implicitly estimating the distribution of incorrect option features for each given context matrix, it is more effective to explicitly estimate the distribution of all RPM image features $z_{ij}$. In this way, when confronted with a specific context matrix, the required incorrect option features can be sampled in a targeted manner from this estimated distribution.
However, accurately estimating the distribution of RPM image features $z_{ij}$ poses significant challenges, and even if such a distribution were successfully estimated, targeted sampling from it would remain highly non-trivial. Therefore, this paper argues that explicitly modeling the distribution of $z_{ij}$ during the process where the DIO’s feature extraction module encodes $x_{i}$ into $z_{ij}$ constitutes an effective approach, since this method eliminates the need for estimation and facilitates targeted sampling through direct distributional control.
This paper has found that the Gaussian Mixture Model (GMM) [36] offers remarkable advantages, as it not only facilitates efficient modeling but also enables targeted sampling. Suppose the features $z_{ij}$ extracted by DIO from RPM images follow a GMM with known numbers of components and parameters. Then, for a given context matrix, after identifying the components corresponding to its correct options within this GMM, we can reasonably regard the components not corresponding to the correct options as the source of incorrect option features. Based on this, we are able to sample comprehensive and diverse incorrect option features from these specific components for DIO. Given these advantages, this section proposes the WORLD method to model the distribution of $z_{ij}$ as a GMM.
VII-B Model the Distribution of RPM Image Features as a Gaussian Mixture Model.
To accurately model the distribution of RPM image features $z_{ij}$ as a GMM, this paper proposes the distribution modeling method WORLD tailored for DIO. Specifically, this paper introduces a trainable vector set $\{U_{m}|m\in[1,M]\}$ and regards it as the mean vectors of a GMM. Using this approach, we construct a predefined GMM with $M$ components. Subsequently, the image features $z_{ij}$ encoded by DIO are projected onto this predefined GMM framework, thereby modeling the distributional characteristics of $z_{ij}$.
The way in which DIO’s related modules extract features from RPM images also only requires minor modifications, so that the distribution of the extracted features $z_{ij}$ can be more easily projected onto the predefined GMM $\{U_{m}|m\in[1,M]\}$. The modifications are shown in Figure 7.
Figure 7: The modified Image Feature Extraction Module.
Specifically, doubling the overall dimension of the ViT within DIO’s image feature extraction module can achieve the desired effect, causing the dimension of the latent features encoded for RPM images to double accordingly. After doubling the dimension, the resulting latent features are evenly split into two parts, denoted as $u_{ij}$ and $b_{ij}$ respectively. Then, the sum of $u_{ij}$ and $b_{ij}$ is taken as the new feature representation $z_{ij}$ for the image $x_{i}$. This process can be expressed by the following formula:
$\{u_{ij},b_{ij}\}=\text{ViT}(x_{i}),\,z_{ij}=u_{ij}+b_{ij}$ (34)
The newly devised derivation method for $z_{ij}$ ensures that, when $b_{ij}$ follows a standard Gaussian distribution $\mathcal{N}(0,1)$, the distribution of $z_{ij}$ explicitly adheres to a Gaussian distribution $\mathcal{N}(u_{ij},1)$. Given this, we merely need to assign the latent features $u_{ij}$ of all observed RPM images onto specific components of the predefined GMM $\{U_{m}|m\in[1,M]\}$ and make the latent image features $b_{ij}$ follow the distribution $N(0,1)$, so as to effectively model the distribution of $z_{ij}$ as a GMM.
The aforementioned method for modeling the distribution of $z_{ij}$ can be implemented by introducing five additional loss function terms.
  1. 1.
    A loss term $\ell_{1}$ aims to encourage the mean of the distribution of $z_{ij}$ to approximate the predefined GMM $\{U_{m}|m\in[1,M]\}$, and its mathematical expression can be formulated as:
    $\ell_{1}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|u_{ij}-\text{gs}[{U}_{\hat{m}}]\|_{2}^{2}$ (35)
    $\text{where}\quad\hat{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{ij}-{U}_{m}\|_{2}^{2}$
    Here, $\text{gs}[\cdot]$ denotes the gradient stopping operation.
  2. 2.
    A loss term $\ell_{2}$ is introduced to enforce $b_{ij}$ to follow a standard normal distribution, expressed as:
    $\ell_{2}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|b_{ij}-\epsilon_{ij}]\|_{2}^{2}$ (36)
    $\text{where}\quad\epsilon_{ij}\sim N(0,1)$
  3. 3.
    A loss term $\ell_{3}$, which is utilized to train the components within the predefined GMM $\{U_{m}|m\in[1,M]\}$, can be expressed as:
    $\ell_{2}=\sum_{i=1}^{16}\sum_{j=1}^{N}\|{U}_{\hat{m}}-\text{gs}[u_{ij}]\|_{2}^{2}$ (37)
    $\text{where}\quad\hat{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{ij}-{U}_{m}\|_{2}^{2}$
  4. 4.
    A mutual information loss term $\ell_{4}$, which facilitates the training of components within the predefined GMM $\{U_{m}|m\in[1,M]\}$, is expressed as:
    $\ell_{4}$ $=\sum_{i=1}^{16}\|x_{i}-\hat{x}_{i}\|_{2}^{2}$ (38)
    where $\hat{x}_{i}=\mathcal{D}\left(\left\{\hat{u}_{ij}+b_{ij}\mid j\in[1,N]\right\}\right),$
    $\hat{u}_{ij}=U_{\hat{m}}-\text{sg}[u_{ij}]+u_{ij},$
    $\hat{m}=\mathop{\mathrm{argmin}}_{m\in[1,M]}\|u_{ij}-U_{m}\|_{2}^{2}.$
    Obviously, the loss term $\ell_{4}$ is formulated to maximize the mutual information between the components within $\{U_{m}|m\in[1,M]\}$ and the latent features $u_{ij}$, via an auxiliary reconstruction task facilitated by the decoder $\mathcal{D}(\cdot)$. The decoder $\mathcal{D}(\cdot)$ is structured as a depth-reversed counterpart of the ViT, transforming latent features into a reconstructed image. The computation of $\hat{u}_{ij}$ ensures the existence of a gradient propagation pathway: $\hat{x}_{i}\rightarrow U_{\hat{m}}\rightarrow\hat{u}_{ij}\rightarrow u_{ij}\rightarrow x_{i}$, thereby guaranteeing the effectiveness of the loss term $\ell_{4}$.
  5. 5.
    A mode collapse prevention loss term $\ell_{5}$ for the latent features $u_{ij}$, defined as:
    $\ell_{5}$ $=\sum_{i=1}^{16}\|x_{i}-\tilde{x}_{i}\|_{2}^{2}$ (39)
    where $\tilde{x}_{i}=\mathcal{D}\left(\left\{\hat{u}_{ij}+\epsilon_{ij}\mid j\in[1,N]\right\}\right),$
    $\epsilon_{ij}\sim N(0,1)$
    $\hat{u}_{ij}=U_{\hat{m}}-\text{sg}[u_{ij}]+u_{ij},$
    $\hat{m}=\mathop{\mathrm{argmin}}_{m\in[1,M]}\|u_{ij}-U_{m}\|_{2}^{2}.$
    It is not hard to notice that during the construction of the loss term $\ell_{4}$, there exists a variable calculation pathway, which can also be referred to as an information pathway, namely: $x_{i}\rightarrow b_{ij}\rightarrow\hat{x}_{i}$. The presence of this pathway may potentially lead to mode collapse in the latent features $u_{ij}$, thereby making it impossible to model the distribution of $z_{ij}$ as a GMM. Therefore, the loss term $\ell_{5}$ comes into being, ensuring that the latent feature $u_{ij}$ cannot be bypassed during the reconstruction process of the RPM image $x_{i}$.
After performing a weighted summation of these five loss terms, we obtain the complete loss function $\ell_{\text{WORLD}}$:
$\ell_{\text{WORLD}}=\ell_{1}+\ell_{2}+0.25\cdot\ell_{3}+\ell_{4}+\ell_{5}$ (40)
By employing $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$ as the loss functions, we can train a DIO model where the distribution of the latent variable $z_{ij}$ is modeled as a GMM. Notably, when training the DIO using $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, the feedforward mechanism undergoes the transformation illustrated in Figure 8.
Figure 8: Forward Feed Process of DIO Model Incorporating $\ell_{\text{WORLD}}$ as an Additional Loss Term.
In figure 8, the latent variable fed into the progressive pattern induction module within DIO is no longer $z_{ij}$, but rather $u_{ij}+\epsilon$. In other words, the forward feed process indicated by Equation (IV-C) is updated as follows:
$\{P_{jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}_{i=\alpha}$
$=\text{PPIM}(\{u_{ij}+\epsilon_{ij},u_{\alpha j}+\epsilon_{\alpha j}|i\in[1,8]\})|_{j=1}^{N}$ (41)
Where $u_{ij}+\epsilon_{ij}$ can be regarded as a sample drawn from the GMM component $U_{\hat{m}}$ that $z_{ij}$ follows. This design enables DIO to interact with a fully modeled GMM rather than deterministic image features, thereby training the DIO to make decisions based on the complete probabilistic structure of the GMM instead of specific feature instantiations.
VII-C Leveraging the Modeled GMM to Provide Incorrect Options for Context Matrices
To steer DIO’s learning objective towards a tighter variational lower bound of the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, we also need to sample sufficient and diverse incorrect option features from the modeled GMM $\{U_{m}|m\in[1,M]\}$ for each given context matrix. Thanks to the inherent properties of GMM, targeted sampling within the set $\{U_{m}|m\in[1,M]\}$ is relatively straightforward. For a given context matrix $\{x_{i}|i\in[1,8]\}$, after identifying the components in the set $\{U_{m}|m\in[1,M]\}$ that correspond to the provided correct option $x_{\alpha}$, samples randomly drawn from the remaining components can be considered, to a certain extent, as features of incorrect options.
Identifying the components corresponding to the correct option, targeted sampling of incorrect option features, and steering the learning objective of DIO can all be achieved by introducing an additional loss function term:
${\ell_{\text{Weighted}}}=$ $-\sum_{j=1}^{N}\log\frac{{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{\tilde{m}}+\epsilon_{\tilde{m}}|i\in[1,8]\})}{|\tau|}}}}}{\sum_{m=1,m\neq\tilde{m}}^{M}{{e^{\frac{\tilde{\text{DIO}}(\{u_{ij}+\epsilon_{ij},U_{m}+\epsilon_{m}|i\in[1,8]\})}{|\tau|}}}}}$ (42)
where $\tilde{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{\alpha j}-{U}_{m}\|_{2}^{2}$
$\epsilon_{m}\sim N(0,1)$
$\tau$ is a trainable smoothing parameter. Given that RPM problems often feature a diverse range of correct options, the rigid approach of designating samples drawn from components that do not correspond to the provided correct option as features of incorrect options carries a potential risk of overgeneralization. To mitigate this sample classification bias, we introduce a smoothing parameter $\tau$. $\tilde{\text{DIO}}$ refers to the DIO without the image feature extraction module, which is capable of processing features into scores. The calculation of $\tilde{m}$ serves to identify the component index in the modeled GMM $\{U_{m}|m\in[1,M]\}$ that corresponds to the feature $z_{\alpha j}$ of the correct option $x_{\alpha}$.
The role of $\ell_{\text{Weighted}}$ is to fine-tune the DIO model. Specifically, when DIO approaches optimal performance under the constraint of the loss function $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, incorporating the additional loss term $\ell_{\text{Weighted}}$ to update the loss function to $\ell_{\text{WORLD}}+\ell_{\text{DIO}}+\ell_{\text{Weighted}}$ creates additional learning capacity and performance improvement potential. This adjustment enables the introduction of supplementary incorrect options during DIO’s late-stage training, thereby steering its learning objective toward a tighter variational lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$, ultimately leading to further performance enhancement.
VIII DIEGO: Methods for Rectifying the Causal Chain in DIO
This study has made two attempts to refine the learning objectives of the DIO model, resulting in the development of two distinct methodology: Brando and WORLD. In this section, we devise a methodology named DIEGO (Direct Intervention for Enforcing Guided Optimal Reasoning), which is designed to rectifying the causal chains modeled by DIO, thereby advancing its reasoning performance.
VIII-A Mutual Information Excels at Problem-Solving Rather Than Reasoning
This paper realizes that setting the maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ as the learning objective of DIO is suboptimal. While mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ directly correlates with accuracy in solving RPM problems, it fails to capture causal relationships inherent in the task design. This creates a critical discrepancy: DIO is optimized for high-accuracy RPM solution generation rather than learning the human-predefined reasoning logic. Both the Brando and WORLD methodologies aim to assist DIO in better maximizing mutual information, yet they fail to address this discrepancy.
Therefore, this paper argues that the causal chain underlying RPM problems should be further modeled. The network architecture of DIO strongly drew reference from the causal chain within RPM problems, with the hope that DIO can autonomously model the causal chain during the training process. However, this can only ensure that DIO is structurally aligned with the causal chain in RPM problems. Whether its latent variables are semantically aligned cannot be guaranteed and remains to be verified.
In order to verify semantic alignment, this paper selects the “Distribute 9” database under the RAVEN problems[19] as the research object. We conduct training on this database using DIO, as well as DIO supplemented with Brando and DIO supplemented with WORLD, respectively. Subsequently, we select 30 instances with strictly identical progressive patterns from this database and verify the pairwise similarity of the progressive pattern representations encoded by DIO for these 30 instances. The calculated $30\times 30$ similarity matrices have been plotted into heatmaps as shown in Figure 9.
Figure 9: Heatmap of Progressive Pattern Representation Extracted from the Distribute 9 under RAVEN Problems.
If DIO accurately models the causal chain embedded in RAVEN problems, then after these 30 instances with the same progressive patterns are encoded into progressive pattern representations by DIO, they should demonstrate high cosine similarity among themselves. However, the three similarity matrices in Figure 9 do not corroborate this expectation. This indicates that, even with auxiliary support from the Brando or WORLD methods, the causal chain modeled by DIO remains systematically misaligned with the ground-truth chains inherent in RPM problems. In addition, this further points out the limitations of using mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ maximization as the learning objective. Given this, this paper proposes that rectifying the causal chain modeled by DIO during its training process will substantially improve performance in solving RPM problems.
VIII-B Rectifying the Causal Chain Modeled by DIO.
In preliminary experiments, we found that DIO exhibits deviations in modeling the segment of the causal chain that involves “abstract attributes $\rightarrow$ progressive patterns of attributes”. This corresponds to a misalign between the results $P_{\alpha{jkl}}$ derived from Formula (IV-C) and human semantics. Therefore, this paper proposes that refining the progressive pattern representations $P_{\alpha{jkl}}$ encoded by DIO would be an effective solution.
This paper has taken note of the metadata provided for each instance in RPM problems, which describes the corresponding instance’s progressive pattern. We argue that incorporating this metadata as a supervisory signal into DIO’s training process to supervise the derivation of $P_{\alpha{jkl}}$ can achieve the objective of rectifying the causal chain modeled by DIO. Thus, we have designed a method named DIEGO. DIEGO involves introducing an additional set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$ for DIO, along with an auxiliary loss function term $\ell_{\text{DIEGO}}$.
Specifically, the DIEGO method first establishes a set of trainable reference vectors $\{V_{f}|f\in[1,F]\}$, where each component represents a distinct metadata modality. Consequently, the reference vector set $\{V_{f}|f\in[1,F]\}$ can serve as a comprehensive dictionary for metadata. The loss term $\ell_{\text{DIEGO}}$ serves to align each progressive pattern representation $P_{\alpha{jkl}}$ extracted by DIO with its corresponding reference vector within $\{V_{f}|f\in[1,F]\}$ according to the accompanying metadata. The loss function $\ell_{\text{DIEGO}}$ can be expressed as follows:
${\ell_{\text{DIEGO}}}=-\sum_{\hat{f}=1}^{F}\sum_{j=1}^{N}\sum_{k=1}^{9}\sum_{l=1}^{2}\text{Meta}_{{\hat{f}l}}\cdot\log\frac{{{e^{({P_{\alpha jkl}}\cdot{V_{\hat{f}}})/|\tau|}}}}{\sum_{f=1}^{F}{{e^{({P_{\alpha jkl}}\cdot{V_{f}})/|\tau|}}}}$ (43)
Here, $\text{Meta}_{\hat{f}l}$ denotes the indication status of the $l$-th metadata entry with respect to the $\hat{f}-$th type of progressive pattern. Meanwhile, $\tau$ represents a trainable positive scaling coefficient. The detailed expression of the loss term $\ell_{\text{DIEGO}}$ demonstrates our methodology for achieving alignment. This method utilizes a cross-entropy loss function with a scaling coefficient to control the cosine similarities between $P_{\alpha_{jkl}}$ and all components in the metadata dictionary $\{V_{f}|f\in[1,F]\}$ according to the indication information from the metadata.
When training the DIO model using $\ell_{\text{DIO}}+\ell_{\text{DIEGO}}$ as the loss function, a solving model that accurately captures the causal chains in RPM problem will be obtained. Previously, some studies have also attempted to treat metadata as additional supervision signals and introduce extra progressive pattern learning tasks into their solving models, aiming to simultaneously enhance the model’s interpretability and reasoning accuracy. However, research led by MRNet[25, 28, 27, 31] indicates that this approach may backfire. Specifically, if metadata is directly incorporated into the training process of the solving models they designed, the model’s reasoning performance will instead decline. This may be logically flawed, but it is indeed an established fact. Nevertheless, this paper argues that the DIO model constructed based on causal chains, coupled with the DIEGO method capable of strategically correcting key segments of the constructed causal chains, holds the key to breaking through this impasse.
IX Solving Generative RPM Problems
Generative RPM problems represent an advanced variant of RPM tasks, where the solving model is required to generate an ideal answer for the context matrix $\{x_{i}|i\in[1,8]\}$ instead of selecting from a predefined bounded option pool.
IX-A DIO And WORLD Solve Generative RPM Problems
The WORLD method not only enhances the performance of DIO but also equips it with the capability to solve generative RPM problems. Specifically, the WORLD method introduces an external GMM $\{U_{m}|m\in[1,M]\}$ and a decoder $\mathcal{D}(\cdot)$ for DIO. Meanwhile, under the WORLD framework, the loss function term $\ell_{\text{Weighted}}$ defined in Formula (42) not only tightens the variational lower bound on the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$ for DIO but also endows DIO with the ability to assign weights to each component in the GMM $\{U_{m}|m\in[1,M]\}$ with respect to the given context matrix features $\{z_{ij}|i\in[1,8],j\in[1,N]\}$. The weights assigned to the component $U_{m}$ for the contextual features $\{z_{ij}|i\in[1,8]\}$ can be calculated as follows:
$S_{{m}{j}}=\tilde{\text{DIO}}(\{z_{i{j}},U_{{m}}|i\in[1,8]\})$ (44)
Based on all computed weights $\{S_{mj}|m\in[1,M],j\in[1,N]\}$, we employ the Gumbel-Softmax trick to draw $N$ components from the GMM $\{U_{m}|m\in[1,M]\}$. This process can be formulated as:
$G_{mj}$ $=-\log(-\log(\text{U}(0,1)))$ (45)
$\tilde{S}_{mj}$ $=S_{mj}+G_{mj}$ (46)
$\hat{m}_{j}$ $=\underset{m}{\text{argmax}}\,\{\tilde{S}_{mj}|m\in[1,M]\}$ (47)
$\hat{U}_{j}$ $=U_{\hat{m}_{j}}$ (48)
These $N$ components $\{\hat{U}_{j}|j\in[1,N]\}$ can be regarded as the distribution of solutions estimated by DIO for the context matrix $\{x_{i}|i\in[1,8]\}$. The decoder $\mathcal{D}$ is capable of transforming the samples from these $N$ components into the generated answer $\hat{x}$:
$\hat{x}=\mathcal{D}(\{U_{\hat{m}_{j}}+\epsilon_{ij}|j\in[1,N]\})$ (49)
$\text{Where}\quad\epsilon_{ij}\sim N(0,1)$
Formulas (34), as well as formulas (44) - (49), indicate that the DIO trained by the WORLD method has the capability to estimate the distribution of solution for the context matrix $\{x_{i}\mid i\in[1,8]\}$ and generate RPM images that can serve as solutions.
IX-B DIEGO Validates the Rationality of Mutual Information Supervision
Verifying the solution accuracy of generative RPM problems has long been challenging. RPM problems often have multiple valid solutions, which means that the generative solving models cannot be evaluated simply by comparing each generated solution with the given correct option. This also makes it challenging for RPM experts to devise a perfect verification algorithm. Most prior work on generative RPM problems has used self-verification, i.e., using their own discriminative models to validate their own generative models.
However, this paper argues that this paradigm of self-verification is inadequate. The cause of the multiple-solution characteristic in RPM problems lies in the fact that solution identification is driven by a progressive pattern. Any RPM image that adheres to the progressive pattern exhibited in the contextual matrices can be considered a valid solution. This implies that algorithms or models used to verify the solution accuracy of generative RPM problems should also be progressive-pattern-driven. The DIEGO approach meets this expectation, as the DIO model trained solely with the loss term $\ell_{\text{DIEGO}}$ will be a progressive-pattern-driven solving model.
Additionally, DIO+WORLD is a strong mutual-information-based inductive model. Using the progressive-pattern-driven model DIEGO to verify DIO+WORLD’s ability to solve generative RPM problems not only avoids the imprecision of symmetric verification but also tests the deeper potential of mutual-information-driven solution models.
X Experiment
All experiments conducted in this paper were programmed in Python, utilizing the PyTorch[37] framework. This paper introduces a comprehensive framework for solving RPM problems through three distinct configurations that integrate a purpose-built deep learning model with specialized training methodologies. The proposed configurations leverage:
  1. 1.
    Solo DIO, which solves RPM problems through a structured causal chain modeling approach without additional optimization strategies. In the DIO model, the hyperparameter $N$ is set to 16, indicating that the ViT within its image feature extraction module employs a $4\times 4$ patch division strategy;
  2. 2.
    Brando-enhanced DIO, which employs trainable hypothetical incorrect options to tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In the Brando framework, the hyperparameter $\beta$ is set to 24, which means the Brando network provides 88 ($\beta-17=8$) constructive hypothetical incorrect options to the DIO in each iteration;
  3. 3.
    WORLD-augmented DIO, utilizing a Gaussian Mixture Model for feature distribution modeling to provide an infinite set of weighted incorrect options for tighting the lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In the WORLD framework, the hyperparameter $M$ is set to 256, equipping the GMM employed for DIO with 256 components. This configuration enables the generation of $M^{N}=256^{16}$ distinct solution distributions, theoretically allowing for the sampling of an infinite diversity of incorrect options.
  4. 4.
    DIEGO-guided DIO, incorporating metadata supervision to directly rectify semantic biases in the causal chain of abstract attribute patterns. In the DIEGO framework, the parameter $F$ is determined by the number of metadata modality in the targeted RPM problem, rather than being arbitrarily set.
This paper conducted experiments on the RAVEN [19], I-RAVEN[34] and PGM [20] datasets, centered around these four configurations. To ensure fair comparison, we uniformly align the optimizer hyperparameters, dataset scales, and batch sizes for optimizing the aforementioned four configurations with those adopted by the current state-of-the-art RS-Tran model [31].
X-A Experiment on RAVEN and I-RAVEN Problems
The performance of our four configurations alongside previous models on the RAVEN and I-RAVEN datasets is summarized in Table I.
TABLE I: Reasoning Accuracies on RAVEN and I-RAVEN.
Test Accuracy(%)
Model Average Center 2 $\times$ 2 Grid 3 $\times$ 3 Grid L-R U-D O-IC O-IG
CoPINet[23] 52.96/22.84 49.45/24.50 61.55/31.10 52.15/25.35 68.10/20.60 65.40/19.85 39.55/19.00 34.55/19.45
PrAE Learner[28] 65.03/77.02 76.50/90.45 78.60/85.35 28.55/45.60 90.05/96.25 90.85/97.35 48.05/63.45 42.60/60.70
SAVIR-T [27] 94.0/98.1 97.8/99.5 94.7/98.1 83.8/93.8 97.8/99.6 98.2/99.1 97.6/99.5 88.0/97.2
MRNet [25] 96.6/- -/- -/- -/- -/- -/- -/- -/-
RS-TRAN[31] 98.4/98.7 99.8/100.0 99.7/99.3 95.4/96.7 99.2/100.0 99.4/99.7 99.9/99.9 95.4/95.4
Triple-CFN[33] 98.9/99.1 100.0/100.0 99.7/99.8 96.2/97.5 99.8/99.9 99.8/99.9 99.9/99.9 97.0/97.3
DIO 98.8/99.3 100.0/100.0 99.6/99.7 96.1/97.4 99.8/99.9 99.8/99.9 99.8/99.8 96.7/97.9
DIO+Brando 99.4/99.5 100.0/100.0 99.8/99.8 98.4/99.0 100.0/100.0 99.9/100.0 100.0/100.0 98.6/98.9
DIO+WORLD 99.5/99.7 100.0/100.0 99.8/99.8 98.8/99.5 99.9/100.0 100.0/100.0 100.0/100.0 99.0/99.5
DIO+DIEGO 99.9/99.9 100.0/100.0 99.9/99.9 99.7/99.8 100.0/100.00 100.0/100.0 100.0/100.0 99.8/99.8
Experimental results recorded in Table I indicate that all four model configurations proposed in this study outperform all existing baselines on both the RAVEN and I-RAVEN datasets. Notably, the DIO+DIEGO configuration achieves a peak accuracy of 99.9%, delivering significant performance gains over prior state-of-the-art models (Triple-CFN[33] and RS-TRAN[31]). Particularly in complex reasoning tasks such as $\times3$ Grid and O-IG, the DIO series models demonstrate distinct advantages, validating the generalizability and robustness of their architectural design for abstract visual reasoning challenges.
More critically, the performance achieved by the DIEGO method has overturned a key experimental conclusion in prior studies[33, 25, 31], which posited that the progressive pattern supervision directly implemented from metadata would inevitably undermine the model’s reasoning capabilities. This logically flawed assertion is obsolete. Furthermore, our experiments demonstrate that when no additional supervision signals are incorporated, the strength of mutual information inferred by the solving model serves as the primary determinant of its problem-solving performance, offering critical architectural insights for future abstract reasoning systems. When progressive pattern descriptions are employed as supervisory signals, the supervision utilization strategy combining causal chain modeling with correction of key chain segments proves highly effective.
X-B Experiment on PGM Problem
The performance comparison among the four model configurations and existing methods on the i.i.d. sub-task “neutral” of the PGM dataset is summarized in Table II.
TABLE II: Reasoning accuracies of models on PGM-neutral.
Model Test Accuracy(%)
CoPINet[23] 56.4
SAVIR-T [27] 91.2
MRNet [25] 94.5
RS-CNN[31] 82.8
RS-TRAN[31] 97.5
Triple-CFN[33] 97.8
DIO 98.1
DIO+Brando 99.0
DIO+WORLD 99.2
DIO+DIEGO 99.6
As shown in Table II, DIO achieves state-of-the-art performance on the PGM-neutral subset of the i.i.d. RPM dataset, a benchmark where prior models exhibited relatively low accuracy. Notably, the integration of mutual information lower bound tightening methods (Brando and WORLD) yields substantial performance improvements for DIO. Furthermore, the DIO+DIEGO attains a peak accuracy of 99.6%, thereby challenging the prevailing hypothesis that metadata supervision necessarily impairs relational reasoning capabilities.
The paper also conducted experiments on the o.o.d. sub-task of PGM, with the results documented in Table III.
TABLE III: Generalization Results on PGM.
Dataset
Model/
Task
Interpolation Extrapolation
Held-out
Attribute
shape-colour
Held-out
Attribute
line-type
Held-out
Triples
Held-out
Pairs of
Triples
Held-out
Attribute
Pairs
MRNet[25]
68.1 19.2 16.9 30.1 25.9 55.3 38.4
RS-Tran[31]
77.2 19.2 12.9 24.7 22.2 43.6 28.4
Triple-CFN[33]
80.4 18.4 12.6 25.2 22.0 44.5 29.2
DIO
82.0 18.8 13.6 26.5 23.3 45.5 29.9
DIO+Brando
83.2 18.2 14.2 26.7 23.6 45.8 30.5
DIO+WORLD
87.8 18.9 15.0 27.6 23.8 46.2 30.9
DIO+DIEGO
92.5 13.8 14.4 27.2 28.8 98.1 98.1
The results recorded in Table III reveal that while mutual information tightening and causal chain modeling yield comparable performance gains for DIO on i.i.d. RPM tasks, the latter demonstrates significantly superior generalization capabilities under o.o.d. RPM conditions. To a certain extent, these results suggest that mutual information tightening remains a highly effective approach for i.i.d. reasoning tasks, despite mutual information being a statistical measure that does not inherently encode causal relationships crucial for reasoning. However, when addressing o.o.d. reasoning challenges, adopting causal chain modeling as the methodological cornerstone proves to be a more robust choice.
X-C Experiment on Generative RAVEN and PGM Problems
While DIO+WORLD can solve discriminative RPM problems with high precision, it is also capable of solving generative RPM problems. Specifically, DIO+WORLD can generate solutions for the context matrices instead of making a selection from a predefined bound option pool. This paper proposes that it is reasonable to use a progressive-pattern-driven discriminative model to verify the accuracy of the generative model in solving generative RPM problems. Therefore, we employ DIO+DIEGO to validate the generative accuracy of DIO+WORLD. Prior to this, we must report DIEGO’s discriminative accuracy for progressive patterns under the supervision of RPM metadata to confirm that DIEGO can effectively verify the solution accuracy for generative RPM problems. DIEGO’s reasoning accuracy for progressive patterns on the RAVEN, I-RAVEN, and PGM datasets is recorded in Table IV.
TABLE IV: DIEGO’s Reasoning Accuracy for Progressive Patterns on the RAVEN, I-RAVEN, and PGM Datasets.
Dataset Sub-task Accuracy(%)
RAVEN center entire:99.9
$\times2$ Grid entire:99.9
$\times3$ Grid entire:99.6
L-R left:99.9 right:99.9
U-D left:99.9 right:99.9
OI-C in:99.9 out:99.9
OIG in:99.8 out:99.9
I-RAVEN center entire:99.9
$\times2$ Grid entire:99.9
$\times3$ Grid entire:99.6
L-R left:99.9 right:99.9
U-D left:99.9 right:99.9
OI-C in:99.9 out:99.9
OIG in:99.8 out:99.9
PGM neutral shape:99.7 line:99.9
Table IV indicates that the DIO+DIEGO demonstrates remarkably high accuracy in progressive pattern reasoning for three categories of RPM problems, making it a suitable validation model for generative RPM tasks. Consequently, we evaluate the DIO+WORLD model using DIO+DIEGO and report its accuracy on generative RAVEN, I-RAVEN and PGM-neutral problems in Table V.
TABLE V: Generation Accuracies of DIO+WORLD
Dataset Sub-task Accuracy (%)
RAVEN/I-RAVEN
Center 91.0/91.2
$\times2$ Grid 85.2/85.0
$\times3$ Grid 81.5/81.8
L-R 87.9/88.5
U-D 87.9/88.4
O-IC 89.0/89.3
O-IG 83.0/84.2
PGM Neutral 84.5
Table V clearly demonstrates that, without relying on any additional supervisory signals, DIO+WORLD has achieved highly competitive results in open-ended generative RPM problems, which strongly confirms the significant research value of the DIO model. In addition, these results also demonstrate that although mutual information does not directly indicate causal relationships, it remains highly valuable in constructing reasoning capabilities and can withstand the tests conducted by models driven by causal chains.
XI Ablation Study
In this section, this paper conducts a series of ablation experiments. The ablation experiments will be conducted focusing on the key parameter settings.
XI-A Ablation Study on DIO
The ablation study on DIO focuses on the design of the network’s feed-forward structure. Within DIO’s progressive pattern induction module, a permutation and combination strategy ($C_{n}^{r}$) is employed to extract progressive patterns from row and column features. This approach aims to model the causal chain details in RPM problems. In this ablation experiment, we will remove the $C_{n}^{r}$ operation. Instead, we will directly input the total of six row-column features for progressive pattern extraction. By doing so, we aim to obscure the structure in DIO that aligns with the causal chain, thereby validating the significance of causal chain modeling as a methodology for network construction. The obscured DIO is denoted as DIO. The reasoning accuracy achieved by DIO on key sub-tasks of RAVEN, I-RAVEN, and PGM is recorded in Table VI.
TABLE VI: Reasoning Accuracies of DIO.
Dataset Sub-task Accuracy (%)
RAVEN/I-RAVEN
$\times3$ Grid 95.6/96.8
O-IG 95.5/95.6
PGM Neutral 97.2
Table VI indicates that constructing RPM solving models based on causal chains is a critically important condition for achieving high reasoning performance.
XI-B Ablation Study on Brando
The ablation study on Brando focuses on the hyperparameter $\beta$. This paper demonstrates that the number of constructive incorrect options provided to DIO directly influences the tightness of the variational lower bound on mutual information derived by DIO. Therefore, it is essential to conduct ablation experiments focusing on the number ($\beta$-17) of hypothetical options provided by the Brando network. The reasoning accuracies of DIO+Brando on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $\beta$ values are recorded in Table VII.
TABLE VII: Reasoning Accuracy of DIO+Brando under Different $\beta$ Values.
Dataset Sub-task Accuracy(%)
$\beta$=24 $\beta$=22 $\beta$=20 $\beta$=18 No $\beta$
RAVEN/
I-RAVEN
$\times3$ Grid 98.4 / 99.0 98.0 / 98.5 97.2 / 98.0 96.8 / 97.8 96.1 / 97.4
O-IG 98.6 / 98.9 98.2 / 98.6 97.6 / 98.3 97.5 / 98.3 96.7 / 97.9
PGM Neutral 99.0 98.8 98.8 98.5 98.1
Table VII indicates that even when only a small number of constructive hypothetical options are introduced into DIO, significant performance improvements can already be achieved for DIO.
XI-C Ablation Study on WORLD
The ablation study on WORLD focuses on the hyperparameter $M$. $M$ indicates the number of components in the GMM constructed by World for RPM image features. The value of $M$ determines the maximum diversity of solution distributions that WORLD can generate for DIO, expressed as $M^{N}$. Therefore, ablation experiments centered around the value of $M$ are warranted. The reasoning accuracies of DIO+WORLD on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $M$ values are recorded in Table VIII.
TABLE VIII: Reasoning Accuracy of DIO+WORLD under Different $M$ Values.
Dataset Sub-task Accuracy(%)
MM=256 MM=128 MM=64 MM=32 No $M$
RAVEN/
I-RAVEN
$\times3$ Grid 98.8 / 99.5 98.5 / 99.1 97.9 / 98.5 96.5 / 97.9 96.1 / 97.4
O-IG 99.0 / 99.5 98.6 / 99.0 98.0 / 98.7 97.1 / 98.0 96.7 / 97.9
PGM Neutral 99.2 99.0 98.5 98.0 98.1
The number of components directly affects the complexity of the GMM. As shown in Table VIII, when the number of GMM components is relatively small ($M=32$), meaning that WORLD models a distribution with lower complexity, it will have a negative impact on the performance of DIO in the PGM task. This study posits that the effectiveness of the WORLD method depends on whether the GMM it constructs can effectively represent the inherent complexity of the natural distribution of features encoded by DIO for RPM images. Specifically, when $M$ is set to 32, this GMM with a low number of components cannot adequately capture the complexity of the natural distribution of PGM image features. Forcing the distribution of PGM image features with higher complexity into this simplified GMM will inevitably lead to a performance degradation of DIO.
WORLD introduces an assumption for DIO, stating that the RPM image features encoded by DIO should conform to a GMM. As a highly versatile distribution model, the GMM theoretically ensures that the WORLD method has a considerable scope of effectiveness. However, the phenomenon where WORLD’s modeling of a GMM with a low number of components causes performance degradation in DIO imposes a crucial prerequisite for the validity of WORLD: the GMM modeled by WORLD must be capable of accommodating the natural complexity of the target problem.
Certainly, in traditional deep learning tasks, underfitting is a prevalent and expected phenomenon that arises when neural networks lack sufficient capacity to handle the complexity of problems. However, the WORLD method makes this phenomenon manageable through its $M$-dependent capacity control. Therefore, particular attention must be paid to the parameter $M$ when applying and adjusting this method.
XII Conclusion
This study has constructed a complete abstract reasoning solution through systematic innovations, with four core innovations forming a progressive technical system that achieves significant progress in both theoretical breakthroughs and empirical performance:
  1. 1.
    Causal Chain Modeling
    • First established a complete causal chain analysis framework of “Image $\rightarrow$ Abstract Attributes $\rightarrow$ Progressive Patterns $\rightarrow$ Pattern Consistency $\rightarrow$ Answer”
    • Provided an explainable causal relationship modeling paradigm for deep learning models.
  2. 2.
    Brando Error Mapping Mechanism
    • Developed a trainable hypothetical incorrect option mapping mechanism to explicitly tighten the variational lower bound of mutual information.
    • Creatively introduced mutual information tightening concepts into abstract reasoning frameworks, thereby opening novel avenues for variational inference methodologies.
  3. 3.
    WORLD Feature Modeling
    • Constructed a Gaussian Mixture Model (GMM) for RPM image feature distributions, enabling infinite and semantically-rich incorrect option sampling
    • Empowering DIO to solve open-ended generative RPM problems.
  4. 4.
    DIEGO Causal Correction Innovation
    • Introduced external metadata to construct supervision signals, precisely correcting semantic biases in the causal chain modeled by DIO
    • Established mappings between model reasoning chains and ground-truth causal chains, aligning internal representations with human cognitive frameworks
[…]
This study’s multi-dimensional innovations not only achieve technical breakthroughs but, more importantly, provide a new paradigm for causal explainability research in deep learning models, holding significant theoretical value and practical implications for advancing artificial intelligence toward higher-order cognitive intelligence.